OpenStudy (anonymous):

Calculus 3 What is the area inside the region bounded by the cardioids r=1+sin(theta) and r=1+cos(theta)?

OpenStudy (amistre64):

have you determined your thetas of intersection .... intervals of concern

OpenStudy (anonymous):

They intersect at pi/4 and the origin.

OpenStudy (amistre64):

somehting like this eh

OpenStudy (amistre64):

due to symmetry, we only need to determine half of it

OpenStudy (anonymous):

Yes, I know what the region looks like, I'm just unsure of how to set up my double integral?

OpenStudy (amistre64):

the hard part is the interval\[\int_{\alpha}^{\beta}\int_{0}^{r(t)}drdt\]

OpenStudy (anonymous):

Oh, would it be\[2\int\limits_{\pi/4}^{\pi}\int\limits_{0}^{1+\cos \theta}r drd \theta \] ??

OpenStudy (amistre64):

im not sure what the theta interval is yet. but yes, that how we would work it

OpenStudy (amistre64):

recall if this is a circle \[\int_{0}^{2\pi}\int_{0}^{R} r drdt\] \[\int_{0}^{2\pi} \frac12R^2dt\] \[\int_{0}^{2\pi}\frac12R^2dt\] \[\frac12R^2~2\pi=\pi R^2\]

OpenStudy (amistre64):

so in this case, we could simply reduce it real quickly to \[\frac22\int_{a}^{b}(1+cos(t))^2~dt\]

OpenStudy (anonymous):

Still no clue??

OpenStudy (amistre64):

still trying to verify the intersection is all

OpenStudy (amistre64):

45 degrees and 225 degrees?

OpenStudy (anonymous):

Yes! Ok just figured it out. \[2\int\limits\limits_{\pi/4}^{5\pi/4}\int\limits\limits_{0}^{1+\cos \theta}r drd \theta\]

OpenStudy (amistre64):